OpenStudy (anonymous):

Jason left a bin outside in his garden to collect rainwater. He notices that 1 over 5 gallon of water fills 2 over 3 of the bin. Write and solve an equation to find the amount of water that will fill the entire bin. Show your work. Explain your answer in words.

OpenStudy (anonymous):

@Michele_Laino @sammixboo

OpenStudy (anonymous):

@Michele_Laino here's the question...

OpenStudy (michele_laino):

I see....

OpenStudy (michele_laino):

I think that in order to fill completely our bin are necessary: \[\frac{ 3 }{ 2 }*\frac{ 1 }{ 5 }\] \[\frac{ 3 }{ 10 }\] of rainwater

OpenStudy (michele_laino):

because if 1/5 gallon of rainwater fills only 2/3 of our bin than, in order to fill completely our bin are necessary (3/2)*(1/5)=3/10

OpenStudy (anonymous):

thx @Michele_Laino

OpenStudy (anonymous):

The price of a video game is $32.99 before tax. Andre bought the video game for 20% off. He then paid 8% sales tax on the discounted price. Part A: How much did Andre pay in sales tax? Show all work and steps in your solution. (5 points) Part B: What is the total amount that Andre paid for the video game? Show all work and steps in your solution. (5 points)

OpenStudy (anonymous):

last one!

OpenStudy (michele_laino):

wait please, we have to write the general eqution as required from your problem!

OpenStudy (anonymous):

oh ok :D

OpenStudy (anonymous):

sorry but i thought that was it lol :D

OpenStudy (michele_laino):

ok! I try to solve your new question....

OpenStudy (anonymous):

ok, is that it for the first one??

OpenStudy (michele_laino):

price before tax due by Andre is: 32.99-(32.99*0.2)=26.39

OpenStudy (anonymous):

oh and is the other question finished??

OpenStudy (michele_laino):

now we have to add taxes of 8%: so cost for videogame is: 26.39+(0.08*26.39)=28.50$ this is part A

OpenStudy (michele_laino):

now I solve for part B...

OpenStudy (michele_laino):

sorry: 28.50 is the part B. Part A is 26.39*0.08=2.11$ of taxes payed by Andre

OpenStudy (anonymous):

so then thats it? your last post is what i have to put??? if it isnt then pls tell me what i have to put in one post

OpenStudy (michele_laino):

your answer is this: Andre paid in sales tax an amount of 26.39*0.08=2.11$ and Andre paid for his videogame an amount of 28.05$

OpenStudy (anonymous):

thx soooooooooo much @Michele_Laino!!!!!!!!! ur the best

OpenStudy (michele_laino):

thank you!!!!! @77777jeannie77777 also you are the best!!!

OpenStudy (anonymous):

thx lol :D

OpenStudy (anonymous):

bye :D

OpenStudy (michele_laino):

thanks again!!!!! bye!!